Solve the system by addition. –3x – 4y = –2
x + 3y = –6

Answer by uma(370)

given -3x - 4y = -2 --------------(1)
x + 3y = -6 -----------------(2)
Direct addition will not eliminate one variable as the coefficients are different.
We first make the coefficients of x equal and then add.
-3x - 4y = -2 -------------(1)
3x + 9y = -18 -------------(2)
Adding the above we get,
5y = - 20
==> 5y/5 = - 20/5
==> y = -4
Plugging this in (2) ,
x + 3(-4) = -6
==> x - 12 = - 6
==> x - 12 + 12 = -6 + 12
==> x = 6
Thus x = 6 and y = - 4

Good Luck!!!